[Hallb] Fwd: [Hps] Mar 27 Day and Swing and Mar 28 Owl CANCELLED

Stepan Stepanyan stepanya at jlab.org
Fri Mar 27 08:49:57 EDT 2015

Hello all,

Due to CHL issues machine restoration has been delayed, at least another 
Will no more this afternoon. In the mean time Hall-B will remain in 
access until we will now more about beam delivery to the hall.

Regards, Stepan

-------- Forwarded Message --------
Subject: 	[Hps] Mar 27 Day and Swing and Mar 28 Owl CANCELLED
Date: 	Fri, 27 Mar 2015 12:46:09 +0000
From: 	Jaros, John A. <john at slac.stanford.edu>
To: 	hps at jlab.org <hps at jlab.org>

Dear HPS,
Day shift and swing shift today and Owl shift tomorrow are cancelled.

The machine has two He liquefier plants. That for the S linac (CHL1) appears operational; that for the N linac (CHL2) has a non-functional
compressor. The problem is being diagnosed today. No time frame for repairs is being discussed until that diagnosis is complete.


Hps mailing list
Hps at jlab.org

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://mailman.jlab.org/pipermail/hallb/attachments/20150327/5bf3ac62/attachment.html 

More information about the Hallb mailing list